A cube with edge of length 4cm is divided into 8 identical cubes. How much greater is the combined surface area of 8 smaller cubes than the surface area of the original cube
Answer by Alan3354(69443)

A cube with edge of length 4cm is divided into 8 identical cubes. How much greater is the combined surface area of 8 smaller cubes than the surface area of the original cube
---------------
Area of 4 cm = 6*4^2 = 96 sq cm
Area of 2 cm cube = 6*2^2 = 24 sq cm
8 of the 2 cm cubes --> 8*24 = 192 sq cm
96 sq cm greater, or 100% greater
--------------
It's 2 times the area = 1 times greater.